# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 9.4. Canonical Correlation Analysis
where $(\xi_{i},\eta_{i})$ are mean 00, variance $1$ Gaussian i.i.d. in $i$ random variables with the squared correlation coefficient of $\xi_{i}$ and $\eta_{i}$ equal to $\theta_{q}$, the asymptotics of $(c_{i}^{2})_{i=1}^{N-1}$ is given to us by Proposition 9.19; $c_{j}=0$ for $j\geq N$. Arguing as in the previous sections, the $q-1$ largest roots of the equation are close to $\lambda_{1},\dots,\lambda_{q-1}$, resulting in (4.6). In order to establish (4.7), we need to investigate (9.72) for $a$ close to $\lambda_{+}=\frac{(\sqrt{\tau_{N}-1}+\sqrt{\tau_{N}-1})^{2}}{\tau_{N}\tau_{M}}$ and locate the root of the equation in the $(\lambda_{q},\lambda_{q-1})$ interval. For this computation, we can approximate $\sum_{i=1}^{S}\xi_{i}^{2}\approx S$, $\sum_{j=1}^{S}\eta_{j}^{2}\approx S$, and $\sum_{i=1}^{S}\xi_{i}\eta_{i}\approx S\sqrt{\theta_{q}}$ because the relative errors in these approximations are of order $N^{-1/2}$, which is smaller than $N^{-1/3}$ scale of our interest. Similarly, the sums $\sum_{j=N}^{M-1}$ can be replaced with their expectations. Hence, up to $O(N^{-1/2})$ error, (9.72) becomes
@@
We change the variables
(9.74) $b=N^{2/3}\frac{a-\lambda_{+}}{\kappa_{1}},\qquad a=\lambda_{+}+N^{-2/3}\kappa_{1}b,$
recall that $\theta^{c}=\frac{1}{\sqrt{(\tau_{N}-1)(\tau_{M}-1)}}$, $\theta_{q}=\frac{1}{\sqrt{(\tau_{N}-1)(\tau_{M}-1)}}+N^{-1/3}\tilde{\theta}$, and apply Theorem 8.20 to (9.73). Arguing exactly as for the factor model in the previous section, we need several applications of the theorem, leading to several functions $\mathcal{G}^{(1)}(b)$, $\mathcal{G}^{(2)}(b)$, …, which are then recombined together. The leading deterministic terms recombine into the same expression as (9.67) evaluated at $a=\lambda_{+}$. Recalling that $m(\lambda_{+})=\mathfrak{m}$, as computed in the proof of Proposition 9.19, the first two lines of (9.73), up to $o(N^{-1/3})$ error, become:
@@
(9.78) $-\tilde{\theta}\kappa_{1}(\tau_{N}-1)^{1/2}(\tau_{M}-1)^{1/2}=2\tfrac{(\tau_{N}-1)^{1/2}(\tau_{M}-1)^{1/2}(1-\lambda_{+})\sqrt{\lambda_{+}}}{\tau_{N}-\tau_{N}\tau_{M}^{-1}-\lambda_{+}(1-\lambda_{+})\mathfrak{m}-\lambda_{+}}\sum_{j=1}^{\infty}\tfrac{\check{\xi}_{j}\bigl{(}(1-\lambda_{+})^{1/2}\eta_{j}-\sqrt{\lambda_{+}}\check{\eta}_{j}\bigr{)}\,}{b-\mathfrak{a}_{j}}\\ -\tfrac{(1-\lambda_{+})\lambda_{+}}{\tau_{N}\tau_{M}^{-1}-1+\lambda_{+}(2-\tau_{N})+\lambda_{+}(1-\lambda_{+})\mathfrak{m})}\sum_{j=1}^{\infty}\frac{\check{\xi}_{j}^{2}}{b-\mathfrak{a}_{j}}-\tfrac{(1-\lambda_{+})}{\tau_{N}\tau_{M}^{-1}+1-\tau_{N}+(1-\lambda_{+})\mathfrak{m}}\sum_{j=1}^{\infty}\frac{((1-\lambda_{+})^{1/2}\eta_{j}-\sqrt{\lambda_{+}}\check{\eta}_{j})^{2}}{b-\mathfrak{a}_{j}}+o(1).$
@@
$\tau_{N}\tau_{M}-\tau_{M}^{2}+(\tau_{M}-2)\bigl{(}\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1}\bigr{)}^{2}=2\sqrt{\tau_{M}-1}(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)(\sqrt{\tau_{M}-1}+\sqrt{\tau_{N}-1})$
Therefore, crucially, (9.79) is the square of a mean 00 Gaussian random variable. The variance of this random variable equals the expectation of (9.79). Recalling that $\check{\xi}_{j}$, $\eta_{j}$, $\check{\eta}_{j}$ are $\mathcal{N}(0,1)$ with covariances
@@
(9.84) $-\tilde{\theta}\kappa_{1}\tau_{M}\tau_{N}^{2}\frac{(\tau_{N}-1)(\tau_{M}-1)}{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2}(\sqrt{\tau_{M}-1}+\sqrt{\tau_{N}-1})^{2}}=\mathcal{G}(b)+o(1).$
@@
(9.85) $-\tilde{\theta}\frac{\tau_{N}^{1/3}(\tau_{N}-1)^{5/6}(\tau_{M}-1)^{5/6}}{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2/3}(\sqrt{\tau_{M}-1}+\sqrt{\tau_{N}-1})^{2/3}}=\mathcal{G}(b)+o(1).$
Using (9.71), recognizing the constant $\kappa_{2}$ in the left-hand side of (9.85), and comparing with Definition 4.2, we arrive at (4.7). ∎

### v2

@@ 9.4. Canonical Correlation Analysis
where $(\xi_{i},\eta_{i})$ are mean 0, variance $1$ Gaussian i.i.d. in $i$ random variables with the squared correlation coefficient of $\xi_{i}$ and $\eta_{i}$ equal to $\theta_{q}$, the asymptotics of $(c_{i}^{2})_{i=1}^{N-1}$ is given to us by Proposition 9.20; $c_{j}=0$ for $j\geq N$. Arguing as in the previous sections, the q1q-1 largest roots of the equation are close to $\lambda_{1},\dots,\lambda_{q-1}$, resulting in (4.6). In order to establish (4.7), we need to investigate (9.72) for $a$ close to $\lambda_{+}=\frac{(\sqrt{\tau_{N}-1}+\sqrt{\tau_{N}-1})^{2}}{\tau_{N}\tau_{M}}$ and locate the root of the equation in the $(\lambda_{q},\lambda_{q-1})$ interval. For this computation, we can approximate $\sum_{i=1}^{S}\xi_{i}^{2}\approx S$, $\sum_{j=1}^{S}\eta_{j}^{2}\approx S$, and $\sum_{i=1}^{S}\xi_{i}\eta_{i}\approx S\sqrt{\theta_{q}}$ because the relative errors in these approximations are of order $N^{-1/2}$, which is smaller than $N^{-1/3}$ scale of our interest. Similarly, the sums $\sum_{j=N}^{M-1}$ can be replaced with their expectations. Hence, up to $O(N^{-1/2})$ error, (9.72) becomes
@@
We change the variables $b=N^{2/3}\frac{a-\lambda_{+}}{\kappa_{1}}$, $a=\lambda_{+}+N^{-2/3}\kappa_{1}b$, recall that $\theta^{c}=\frac{1}{\sqrt{(\tau_{N}-1)(\tau_{M}-1)}}$, $\theta_{q}=\frac{1}{\sqrt{(\tau_{N}-1)(\tau_{M}-1)}}+N^{-1/3}\tilde{\theta}$, and apply Theorem 8.20 to (9.73). Arguing exactly as for the factor model in the previous section, we need several applications of the theorem, leading to several functions $\mathcal{G}^{(1)}(b)$, $\mathcal{G}^{(2)}(b)$, …, which are then recombined together. The leading deterministic terms recombine into the same expression as (9.67) evaluated at $a=\lambda_{+}$. Recalling that $m(\lambda_{+})=\mathfrak{m}$, as computed in the proof of Proposition 9.20, the first two lines of (9.73), up to $o(N^{-1/3})$ error, become:
@@
(9.77) $\tilde{\theta}\kappa_{1}(\tau_{N}-1)^{1/2}(\tau_{M}-1)^{1/2}=2\tfrac{(\tau_{N}-1)^{1/2}(\tau_{M}-1)^{1/2}(1-\lambda_{+})\sqrt{\lambda_{+}}}{\tau_{N}-\tau_{N}\tau_{M}^{-1}-\lambda_{+}(1-\lambda_{+})\mathfrak{m}-\lambda_{+}}\sum_{j=1}^{\infty}\tfrac{\check{\xi}_{j}\bigl((1-\lambda_{+})^{1/2}\eta_{j}-\sqrt{\lambda_{+}}\check{\eta}_{j}\bigr)\,}{b-\mathfrak{a}_{j}}\\ -\tfrac{(1-\lambda_{+})\lambda_{+}}{\tau_{N}\tau_{M}^{-1}-1+\lambda_{+}(2-\tau_{N})+\lambda_{+}(1-\lambda_{+})\mathfrak{m})}\sum_{j=1}^{\infty}\frac{\check{\xi}_{j}^{2}}{b-\mathfrak{a}_{j}}-\tfrac{(1-\lambda_{+})}{\tau_{N}\tau_{M}^{-1}+1-\tau_{N}+(1-\lambda_{+})\mathfrak{m}}\sum_{j=1}^{\infty}\frac{((1-\lambda_{+})^{1/2}\eta_{j}-\sqrt{\lambda_{+}}\check{\eta}_{j})^{2}}{b-\mathfrak{a}_{j}}+o(1).$
@@
$\tau_{N}\tau_{M}-\tau_{M}^{2}+(\tau_{M}-2)\bigl(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1}\bigr)^{2}=2\sqrt{\tau_{M}-1}(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)(\sqrt{\tau_{M}-1}+\sqrt{\tau_{N}-1})$ .
Therefore, crucially, (9.78) is the square of a mean 0 Gaussian random variable. The variance of this random variable equals the expectation of (9.78). Recalling that $\check{\xi}_{j}$, $\eta_{j}$, $\check{\eta}_{j}$ are $\mathcal{N}(0,1)$ with covariances
@@
(9.83) $\tilde{\theta}\kappa_{1}\tau_{M}\tau_{N}^{2}\frac{(\tau_{N}-1)(\tau_{M}-1)}{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2}(\sqrt{\tau_{M}-1}+\sqrt{\tau_{N}-1})^{2}}=\mathcal{G}(b)+o(1).$
@@
(9.84) $\tilde{\theta}\frac{\tau_{N}^{1/3}(\tau_{N}-1)^{5/6}(\tau_{M}-1)^{5/6}}{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2/3}(\sqrt{\tau_{M}-1}+\sqrt{\tau_{N}-1})^{2/3}}=\mathcal{G}(b)+o(1).$
Using (9.71), recognizing $\kappa_{2}$ and comparing with Definition 4.2, we arrive at (4.7). ∎